# Data-driven rainfall prediction at a regional scale: a case study with Ghana

### v2

@@ 5.4 Input variables selection
We run the Gibbs sampler described in Section 4.3 for $1,000$ epochs, where each iteration is a full sweep through the $K$ variables. We use $\sigma^{2}=0.01$, and $r=3.76$. We discard the first 950 epochs as burn-in. For each input variable $j$, we take the average of $\delta_{j}$ along the MCMC output as an estimate of the posterior probability of $\delta_{j}$. Figure 9 shows the $30$ largest posterior probabilities. The results show that the most important predictive variable is the spatio-temporal variable Cos defined in (1). This is hardly surprising since rainfall in Ghana is known to have a strong seasonality which varies with latitude. Evaporation drives rainfall, and our method indeed highlights specific humidity ($q925$), relative humidity ($r950$) and total column water vapor ($tcwv$) as important input variables. The variable wind ($u300$) also appears important. The wind variable $u300$ is possibly related to the African Easterly Jet (AEJ), which plays a fundamental in the West African monsoon. However, additional research is needed on this connection, since the bulk of the AEJ is typically located much lower, at around $700$hPa. Our methodology also highlights several convection-related variables: convective inhibition ($cin$), K-index ($kx$) and the convective available potential energy ($cape$) as key input variables. This is consistent, since a large part of the rainfall in Ghana is convection-driven. One important limitation of our methodology is that it does not offer much insight on the physical process(es) by which these variables affect rainfall. More research is needed on this point.
@@ 6 Conclusion
This research dealt with the challenging problem of predicting rainfall in tropical regions. Although we have focused on Ghana, our methodology can be readily applied to other regions. Our proposed data-driven model is a U-Net convolutional neural network model (CNN) that we trained to predict 24h rainfall at 12h and 30h lead-time using data from ERA5 and GPM-IMERG. As illustrated, the 12h lead-time prediction model shows significant improvement over a numerical weather prediction (NWP) model developed by the ECMWF as obtained through TIGGE. Our investigation also found that a strategy that mixes our U-Net model and NWP further enhanced forecasting performance. We also developed a novel methodology that enhanced the interpretability of our model, by identifying the input variables that best capture the performance of the model.
@@
Our work comes with several limitations, but also opens up several avenues for further research that we hope to tackle in the future. Although we have identified a small list of variables that appear to play an important role in the forecasts, the precise physical processes by which these variables impact rainfall remain largely unclear. There is a growing literature that combines statistics, machine learning and partial differential equations and aims to discover new physical laws from data (Raissi et al., 2019; Chen et al., 2020). Extending these methods to atmospheric sciences can provide a powerful empirical framework to identifying the physical processes involved in the present study.
Another major limitation of our work is the lack of probabilistic forecast and uncertainty quantification (UQ). Fully addressing this issue in a way that captures correctly both the aleatoric and epistemic uncertainties requires further methodological advances than what is currently available in the literature. We refer the reader to (Haynes et al., 2023) for a good review and illustration of the current state-of-the-art on UQ in environmental sciences. Clearly, some degree of uncertainty quantity can be easily achieved by training an ensemble of our U-Net models. However, the practical reliability of these methods remains limited. Particularly, it is unclear whether these methods can reliably capture the aleatoric uncertainty (see Haynes et al. (2023) for further discussion). Another interesting methodology for UQ from single model outputs is the EasyUQ of (Walz et al., 2024b). However, although very easy to implement, the methodology relies on a fundamental monotonicity assumption that in general is untestable. There is a burgeoning literature that aims to develop rigorous Bayesian methodology for training deep learning models (Wilson and Izmailov, 2020; Atchade and Wang, 2023), with suitable built-in UQ. However the implementation cost of these algorithms for the moment seems prohibitive. Overall, given the state of the literature, and despite the fundamental importance of UQ in predictive modeling, we have chosen to defer this issue to future research.
@@
Another notable shortcoming of our method is its poor performance in predicting heavy and extreme rainfall events, driven mainly by the lack of sufficient tail event data. Models for better prediction of tail-events in statistics are often built using quantile regression. However, quantile regression at extreme quantile levels suffer from the same lack of tail data as in this study. New methodology are thus needed to address this issue, perhaps along the emerging literature of extreme quantile regression that merges classical quantile regression and extreme value theory (Velthoen et al. (2023); Pasche and Engelke (2024)).
@@ Acknowledgments
We are grateful to Nils Cazemier for assistance with the GPM-IMERG data product. We are grateful to Tilmann Gneiting for helpful discussions, and for pointing out several relevant papers. We also acknowledge the financial support of the National Science Foundation grant DMS-2210664.

### v3

@@ 5.4 Input variables selection
We run the Gibbs sampler described in Section 4.3 for $1,000$ epochs, where each iteration is a full sweep through the $K$ variables. We use $\sigma^{2}=0.01$, and $r=3.76$. We discard the first 950 epochs as burn-in. For each input variable $j$, we take the average of $\delta_{j}$ along the MCMC output as an estimate of the posterior probability of $\delta_{j}$. Figure 8 shows the $30$ largest posterior probabilities. The results show that the most important predictive variable is the spatio-temporal variable $\rm{COS}$ defined in (1). This is hardly surprising since rainfall in Ghana is known to have a strong seasonality which varies with latitude. Evaporation drives rainfall, and our method indeed highlights specific humidity ($q925$), relative humidity ($r950$), and total column water vapor ($tcwv$) as important input variables. The variable wind ($u300$) also appears important. The wind variable $u300$ is possibly related to the Tropical Easterly Jet (TEJ), which plays a fundamental role in the West African monsoon (Nicholson, 2013)]. Our methodology also highlights several convection-related variables: convective inhibition ($cin$), K-index ($kx$) and the convective available potential energy ($cape$) as key input variables. This is realistic, since a large part of the rainfall in Ghana is convection-driven. One important limitation of our methodology is that it does not offer much insight on the physical process(es) by which these variables affect rainfall. More research is needed on this point.
@@ 6 Conclusion
This research dealt with the challenging problem of predicting rainfall in tropical regions. Although we have focused on Ghana, our methodology can be readily applied to other regions. Our proposed data-driven model is a U-Net convolutional neural network model (CNN) that we trained to predict 24h rainfall at 12h and 18h lead-time using data from ERA5 and GPM-IMERG. We find that the 12h lead-time prediction model shows improvement in terms of precision and accuracy over a numerical weather prediction (NWP) model developed by the ECMWF as obtained through TIGGE. We also explore with mixed results a hybrid approach that combines the U-Net model with the ECMWF forecasts. In particular, while the hybrid model achieves the lowest MAE among all models and generally outperforms the NWP, it performs worst than the stand-alone U-Net models in terms of CRPS and the ability to detect large rainfall events. These results suggest that more sophisticated methods are needed to fully leverage the potential of the hybrid approach. With the dazzling rise of deep learning in weather forecasting, this issue stands as an important direction for future research. Finally, we also developed a novel methodology that enhanced the interpretability of our model by identifying the input variables that best capture the performance of the model.
@@
Our work comes with several limitations, but also opens up several avenues for further research that we hope to tackle in the future. Although we have identified a small list of variables that appear to play an important role in the forecasts, the precise physical processes by which these variables impact rainfall remain largely unclear. There is a growing literature that combines statistics, machine learning, and partial differential equations and aims to discover new physical laws from data (Raissi et al., 2019; Chen et al., 2020). Extending these methods to atmospheric sciences can provide a powerful empirical framework for identifying the physical processes involved in the present study.
@@
Another notable shortcoming of our method is its poor performance in predicting heavy and extreme rainfall events, driven mainly by the lack of sufficient tail event data. Models for better prediction of tail-events in statistics are often built using quantile regression. However, quantile regression at extreme quantile levels suffer from the same lack of tail data as in this study. New methodologies are thus needed to address this issue, perhaps along the emerging literature of extreme quantile regression that merges classical quantile regression and extreme value theory (Velthoen et al., 2023; Pasche and Engelke, 2024). Expanding the area of interest and/or turning to gauge datasets such as the precipitation products of the Global Precipitation Climatology Centre (Schamm et al., 2014) may also prove useful in addressing this issue.
@@ Acknowledgments
We are grateful to Nils Cazemier for assistance with the GPM-IMERG data product. We are grateful to Tilmann Gneiting for helpful discussions, and for pointing out several relevant papers. We also acknowledge the financial support of the National Science Foundation grant DMS-2210664. The authors also declare no conflicts of interest related to this study.